OpenStudy (anonymous):

How can I find the sum of the first 365 numbers in a sequence? The specific sequence I was given is 3, 6, 9, 12, 15, ..., increasing by 3 each time. It represents somebody saving pennies every day for a year, putting 3 in the first day, 6 in the second, 9 in the third, and so on. I need to know how much money they will have by the end of a year (not a leap year).

OpenStudy (anonymous):

I've only ever done these problems by actually going through and finding all the terms and manually adding them together, but 365 is a bit much, there's gotta be a faster way to do it

OpenStudy (campbell_st):

this is an arithmetic sequence. the formula for the sum of the sequence is \[S_{n} = \frac{n}{2}[2a + (n - 1) \times d]\] n = number of terms in the sequence, in your question n = 365 a = 1st term, in your sequence a = 3 d = common difference, or what is is always increasing by d = 3 substitute these values into the formula above and you'll get the sum

OpenStudy (anonymous):

So it would be 365/2 [2(3)+(365-1)*3

OpenStudy (campbell_st):

thats it

OpenStudy (anonymous):

I think I got confused when solving it... I got 202020

OpenStudy (anonymous):

Is that what I should have gotten?

OpenStudy (anonymous):

And if I convert that from pennies to dollars it would become $2,020.20?

OpenStudy (campbell_st):

I got 200385

OpenStudy (anonymous):

I just calculated it again and got 200,385

OpenStudy (anonymous):

Oh lol

OpenStudy (campbell_st):

and the conversion is correct...

OpenStudy (anonymous):

Alright so it's $2,003.85 (:

OpenStudy (anonymous):

Thank you!